### v1

# SEKE: Specialised Experts for Keyword Extraction

Abstract
Keyword extraction involves identifying the most descriptive words in a document, allowing automatic categorisation and summarisation of large quantities of diverse textual data. Relying on the insight that real-world keyword detection often requires handling of diverse content, we propose a novel supervised keyword extraction approach based on the mixture of experts (MoE) technique. MoE uses a learnable routing sub-network to direct information to specialised experts, allowing them to specialize in distinct regions of the input space. SEKE, a mixture of Specialised Experts for supervised Keyword Extraction, uses DeBERTa as the backbone model and builds on the MoE framework, where experts attend to each token, by integrating it with a recurrent neural network (RNN), to allow successful extraction even on smaller corpora, where specialisation is harder due to lack of training data. The MoE framework also provides an insight into inner workings of individual experts, enhancing the explainability of the approach. We benchmark SEKE on multiple English datasets, achieving state-of-the-art performance compared to strong supervised and unsupervised baselines. Our analysis reveals that depending on data size and type, experts specialize in distinct syntactic and semantic components, such as punctuation, stopwords, parts-of-speech, or named entities. Code is available at https://github.com/matejMartinc/SEKE_keyword_extraction.
1 Introduction
Keyword extraction is crucial for efficiently summarizing, indexing, and retrieving relevant information from large textual corpora, making it one of the most fundamental NLP tasks Song et al. (2023). Automatically extracting keywords is defined as the automatic extraction of words that represent crucial semantic aspects of the text. While the first automated solutions have been proposed more than two decades ago (Tumey, 1999; Mihalcea and Tarau, 2004), the task is far from solved, and the scientific community is still actively trying to improve the performance of the keyword extractors, proposing several new algorithms in recent years (see Section 2 for details).
Recently, the so-called foundation models have completely revolutionised the way natural language processing (NLP) is done, and the development of large language models (LLMs) has enabled task-agnostic architectures to solve downstream NLP tasks in a zero-shot fashion without the need for labelled data Brown et al. (2020). These breakthroughs have been enabled by the integration of novel training regimes and architectural components, such as Low-Rank Adaptation (LoRA) Hu et al. (2021) and Mixture of Experts (MoE) Jacobs et al. (1991), which make models more efficient, while requiring less computational resources.
While these components have been successfully employed in an unsupervised language model setting with abundant data, there are less studies exploring their effectiveness in a supervised setting with much fewer resources. The main reason for this is that these components were developed with a specific objective of improving the scalability of the models Shazeer et al. (2017) and therefore their effect on the performance of the models is somewhat neglected or measured only indirectly, that is, studies focus on how the use of these components allows for greater scalability, which in turn leads to performance gains.
In contrast, in this study, we focus on the performance improvements obtained from these components without increasing the size of the backbone model. More specifically, we are interested in exploring the question of whether MoE can improve the performance of the model on a specific downstream task, in our case keyword extraction. The general hypothesis we want to test is whether it is beneficial for keyword extraction to allow specific parts of the network to specialize for specific types of tokens. More specifically, we explore if introducing a gating network that assigns different parts of the sequence to specialised layers, which would only attend to specific subsequences and tokens according to their semantic and grammatical role, could improve the current state-of-the-art (SOTA) in the field. In addition, we explore what type of specialisation different experts undertake and how much data is required for a successful specialisation. For these reasons, we test our approach on several datasets with different amounts of training data available, to determine the data threshold that still allows the convergence of expert layers. Finally, we examine whether synergy can be achieved between MoE and recurrent neural network (RNN) that would lower this threshold.
The contributions of the paper are as follows.
  • We propose a novel token classification head architecture (see Figure 1) that combines MoE and RNN. The approach outperforms SOTA-supervised sequence-labelling models on most keyword extraction datasets. As far as we know, this is the first research that studies the usage of MoE in a supervised sequence-labelling scenario.
  • We demonstrate that experts specialize in different syntactic and semantic structures, ranging from punctuation characters and part-of-speech tags to specific words and named entities, depending on the domain they are trained on and the size of the training data.
  • We show that adding MoE layers during fine-tuning generally has a positive effect on the model’s performance no matter the dataset, but is somewhat dependent on the amount of data available. In addition, we show that the data requirements can be lowered by introducing additional recurrent layers.
Figure 1: The general architecture given the input.
2 Related Work
2.1 Mixture of Experts
The idea for Mixture of experts (MoE) comes from the study by Jacobs et al. (1991). Like ensemble methods, the concept involved a supervised procedure for a system composed of separate networks, each responsible for a different subset of training cases. Each network, or expert, specializes in a distinct region of the input space. The choice of expert is managed by a gating network, which determines the weights for each expert. Both the experts and the gating network are trained during the training process.
More recently, the study by Eigen et al. (2013) proposed integrating MoE as components within deep neural networks. This allowed MoE to function as layers in a multilayer network, enabling models to be both large and efficient, which led to the employment of the MoE method in the NLP setting. The study by Shazeer et al. (2017) proposed a 137-billion-parameter LSTM model (the prevalent NLP architecture for language modelling at the time) by introducing sparsity, which allowed for very fast inference even at large scales. This work, which focused on translation, encountered several challenges such as high communication costs and training instabilities. After the success of the Mixtral 8x7B Jiang et al. (2024), MoE have been used in several LLMs because they enable more efficient pre-training with fewer computing resources, allowing greater scalability in terms of model size and data set size. The studies show that a MoE model achieves the same quality as its dense counterpart much faster during pre-training, leading to its usage in training multi-trillion parameter models, such as the 1.6T parameters Switch Transformers Fedus et al. (2022).
Although several studies showed that MoE are an efficient strategy to achieve greater scalability, very few studies (if any) focused on performance gains that could be obtained by this method. To our knowledge, no study to date explored how MoE can benefit supervised keyword extraction (or in fact any sequence labelling) tasks.
2.2 Keyword Extraction
Contemporary studies on keyword extraction treat it either as a text generation or a sequence labelling task. Yuan et al. (2020) proposed an encoder-decoder RNN architecture featuring two mechanisms, semantic coverage and orthogonal regularisation, which treated keyword extraction as text generation. These mechanisms ensure that the overall inner representation of a generated keyword sequence is semantically aligned with the overall meaning of the source text. In contrast, several recent sequence labelling approaches tackle the task with transformers. Sahrawat et al. (2020) tested several transformer architectures, namely BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and GPT-2, and two distinct sequence labelling heads, a bidirectional Long short-term memory head (BiLSTM) and a BiLSTM head with an additional Conditional random fields layer (BiLSTM-CRF). The TNT-KID approach Martinc et al. (2022) on the other hand adapted the transformer architecture specifically for keyword extraction tasks by re-parameterizing the attention mechanism to focus more on positional information. Finally, Sun et al. (2021) proposed JointKPE, an open domain keyphrase extraction (KPE) architecture also based on pre-trained transformer language models. It ranks keyphrases by assessing their informativeness across the entire document and is simultaneously trained on a keyphrase chunking task to ensure the phraseness of the keyphrase candidates.
Just recently, Luo et al. (2023) proposed Diff-KPE, a text diffusion approach to generate enhanced keyphrase representations utilizing the supervised Variational Information Bottleneck (VIB). Diff-KPE generates keyphrase embeddings conditioned on the entire document using BERT and then injects the generated keyphrase embeddings into each phrase representation. By optimizing the ranking network and VIB simultaneously, they manage to rank each candidate phrase by utilizing both the information of keyphrases and the entire document.
Keyword extraction can also be tackled in an unsupervised manner. In general, these approaches can be divided into four main categories, namely statistical, graph-based, embeddings-based, and language model-based methods. Statistical methods, such as YAKE Campos et al. (2020), use the statistical characteristics of texts to capture keywords. Graph-based methods, such as TextRank Mihalcea and Tarau (2004) and RaKUn Škrlj et al. (2019) build graphs to rank words based on their position in the graph. Embedding-based methods, such as Key2Vec Mahata et al. (2018), employ semantic information from distributed word and sentence representations for keyword extraction.
Notably, large language model-based methods tend to be highly effective, driven by the rise of LLMs, such as ChatGPT Brown et al. (2020). These models have been employed in several recent studies Martínez-Cruz et al. (2023); Luo et al. (2023), demonstrating promising results due to their zero-shot capabilities, which allow them to perform tasks without task-specific fine-tuning Kojima et al. (2022). However, Zhang et al. (2024) has recently highlighted the theoretical drawbacks of autoregressively trained LLM models for semantic classification tasks, attributing these limitations to their fixed token positioning, which restricts inter-sample connections. In contrast, masked language models, such as BERT Devlin et al. (2019), leverage flexible token targeting to overcome these issues. Building on these findings, we select the DeBERTa He et al. (2020) model as a strong backbone.
3 Methods
3.1 Architecture
Our approach (visualised in Figure 1) is based on fine-tuning a pre-trained transformer architecture, in our case DeBERTa He et al. (2020), with a specialised token classification head. The architecture was chosen since it showcased a SOTA performance on several downstream NLP tasks, among them also sequence-labelling tasks such as NER Shon et al. (2022). Additionally, we hypothesised that the model is especially appropriate for keyword extraction due to its novel disentangled attention mechanism. More specifically, in the DeBERTa model, each word is represented using two vectors that encode its content and position separately, and the attention weights among words are computed using disentangled matrices on their contents and relative positions, respectively. We assumed that this would allow the model to better distinguish between the positional and semantic/lexical information and therefore make it possible to assign attention to some tokens purely based on their position in the text, improving the overall performance. This hypothesis is based on previous study by Martinc et al. (2022) showing that token position is especially important in the keyword identification task due to the disproportionate distribution of keywords in the document, which is skewed towards the beginning of the document (i.e., more keywords appear at the beginning of the document). By considering the importance of positional information, they fed the positional encoding to the attention mechanism directly, subsequently managing to improve the performance.
The main novelty of our approach is a customised token classification head containing several components. The output logits from the DeBERTa backbone are first fed to the gate network or router that determines which tokens in the input sequence are sent to which expert. More specifically, in the setting with n experts, the gate network (G) decides which experts (E) receive a part of the input:
$y=\sum_{i=1}^{n}G(x)_{i}E_{i}(x)$
We model the routing as a weighted multiplication, where hypothetically all experts could contribute to processing each part of the input. Nevertheless, the additional top_k parameter ensures that only top k experts contribute to the respective part of the input, and other experts are excluded from the computational graph, hence saving computational resources. Following Shazeer et al. (2017), we employ the so-called Noisy Top-k Gating, which introduces some (tunable) noise and then keeps the top k values. Noise is introduced for load balancing, i.e., to prevent the gating network from converging in a way that the same few experts are activated for most of the tokens, which would make training inefficient. The final weight distribution is obtained by feeding the input ($x$) through a dense layer ($W_{g}$) and adding some random noise in the following way:
$H(x)_{i}=(x\cdot W_{g})_{i}+\textrm{RND}\cdot\textrm{Softplus}((x\cdot W_{noise})_{i})$
Here, RND represents a weight vector of random numbers from a normal distribution with mean 0 and variance 1. After that, we filter out only the top k experts out of v:
$KeepTopK(v,k)_{i}=\{v_{i}\textrm{ if }v_{i}\textrm{ in top k, else}-\inf\}$
We finally apply the softmax to filtered experts:
$G(x)=\textrm{Softmax}(\textrm{KeepTopK}(H(x),k))$
Note that the router is composed of learnable parameters and is fine-tuned at the same time as the rest of the network. On the other hand, each expert is represented as a neural network consisting of three sequential dense layers $W_{g}$ with additional activation and dropout layers. More specifically, the expert is represented as:
$E(x)=\textrm{Dropout}(\textrm{ReLU}(x\cdot W_{g})\cdot W_{g})\cdot W_{g}$
The outputs of different experts, which attend to specific tokens assigned to them by the routing network, are first reassembled into a single sequence representation. Then, a two-layer randomly initialised encoder, consisting of dropout and two RNN layers, is added (with element-wise summation) to this sequence. The initial motivation behind this adaptation is connected with findings from the related work which suggest that recurrent layers are quite successful at modelling the positional importance of tokens in the keyword detection task and by the study of Martinc et al. (2022), who reported good results when an RNN classifier and contextual embeddings generated by transformers were employed for keyword detection.
[…]
We conduct additional post-processing, namely, the identified keyphrases are extracted from the labelled sequence produced by the model, and we conduct additional filtering, where we remove keyphrases containing punctuation (except dashes and apostrophes), and deduplication, where we remove duplicate keyphrases. The final output list contains a maximum of 10 keyphrases per input document.
4 Experimental Setting
4.1 Datasets
We conduct experiments on selected benchmark datasets used in the SOTA papers to which we compare our approach. More specifically, we experiment on 6 datasets from the related work Meng et al. (2017); Martinc et al. (2022); Xiong et al. (2019), covering three domains, news, scientific articles, and web pages.
We present the datasets in Table 1, which contains information about the domain, the number of train, validation, and test documents, as well as average number of keywords per document. There is notable variability in a number of keywords per document, with some datasets requiring extraction of twice (or even three times) as many keywords as others.
Dataset Domain Train Valid Test Train Valid Test
#Docs #Docs #KW #KW #KW
per doc per doc per doc
KP20k CS 530,000 20,000 20,000 3.30 3.31 3.32
Inspec CS / 1,500 500 / 7.21 7.51
Krapivin CS / 1,844 460 / 2.87 3.22
KPTimes News 259,923 10,000 10,000 2.34 2.31 2.35
JPTimes News - - 10,000 / / 3.84
openKP Web 134,894 6,616 6,614 2.00 1.90 1.92
Table 1: Statistics for keyword extraction datasets. This table presents the number of documents in the train, validation, and test sets for each dataset. The Domain column indicates the source of the datasets (computer science (CS), news, or web), while #KW per doc shows the average number of keywords per document.
4.2 Experiments
We test different settings to determine the effect that the proposed architectural additions have on the performance of the model. We are interested in the specific contributions of the MoE and RNN layers to the performance of the model (either separately or together); therefore, we compare the settings containing these components to the baseline DeBERTa model with just a (usual) feed-forward token classification head.
We employ the DeBERTa model [^1] More specifically, the “deberta-v3-base” version available at https://huggingface.co/microsoft/deberta-v3-base. and fine-tune it using Low-Rank Adaptation (LoRA) Hu et al. (2021). We freeze all layers in the backbone model and train only low-rank perturbations to query and value weight matrices in the model. Additionally, we train all the matrices in the custom token classification head (i.e., MoE, RNN, and dense classification layers are randomly initialised and fine-tuned). We use the same hyperparameter setting across all datasets, namely, the following values were used during fine-tuning: learning rate of 2e-4, LoRA intrinsic rank of 16, LoRA alpha of 16, LoRA dropout of 0.1, and sequence length of 256 for all tasks.
In order to determine how does training dataset size effect our approach, we employ two distinct training scenarios. In the validation set training scenario, the DeBERTa model was fine-tuned just on each dataset’s validation sets, which were randomly split into 80 percent of documents used for fine-tuning and 20 percent of documents used for hyperparameter optimisation and test set model selection. We fine-tune the models for up to 20 epochs and employ early stopping. Note that the JPTimes dataset with no available predefined validation-test splits is only used for testing. Here, the model was fine-tuned on the KPTimes-valid dataset.
In the full training scenario, we fine-tune the models on all available data, same as in related work Meng et al. (2017); Martinc et al. (2022). More specifically, before fine-tuning the model on three relatively small validation datasets containing scientific papers (KP20k-valid, Inspec-valid, and Krapivin-valid), we first “pre-train” the model (using the same token classification objective as in the validation dataset fine-tuning stage) for 20 epochs on the large KP20k-train dataset also containing scientific articles. We do the same for news datasets, i.e., before fine-tuning the model on the KPTimes-valid dataset, we “pre-train” it on the large KPTimes-train dataset. For the JPTimes dataset with no available predefined validation-test splits, the keyword detection model was “pre-trained” on KPTimes-train and fine-tuned on the KPTimes-valid dataset. For the openKP dataset, the model is “pre-trained” on the openKP-train dataset.
To assess the performance of our models and compare the results of our keyword extraction approach to other SOTA methods, we adopt the same evaluation methodology as in Meng et al. (2017); Martinc et al. (2022); Luo et al. (2023). We use the same test splits for the sake of comparability between our approach and others, and measure F1@$k$ with $k\in\{1,3,5\}$ on the openKP dataset and $k\in\{5,10\}$ on other datasets. To determine the exact match, we first lowercase the candidate keywords and the gold standard keywords, and then apply Porter Stemmer Porter (1980) to both.
KP20k Inspec Krapivin KPTimes JPTimes openKP
F1@5 F1@10 F1@5 F1@10 F1@5 F1@10 F1@5 F1@10 F1@5 F1@10 F1@1 F1@3 F1@5
Unsupervised baselines
TF-IDF 7.2 9.4 16.0 24.4 6.7 9.3 17.9 15.1 26.6 22.9 19.6 22.3 19.6
TextRank 18.0 15.0 28.6 33.9 18.5 16.0 2.2 3.0 1.2 2.6 5.4 7.6 7.9
ChatGPT 23.2 10.8 35.2 33.9 12.8 11.7 27.9 / / / 20.8 20.4 16.6
ChatGLM2-6b / / 25.1 30.1 / / / / / / 16.0 11.0 8.6
Supervised baselines
CopyRNN 31.7 27.3 24.4 28.9 30.5 26.6 40.6 39.3 25.6 24.6 21.7 23.7 21.0
CatSeqD 34.8 29.8 27.6 33.3 32.5 28.5 42.4 42.4 23.8 23.8 / / /
GPT-2 27.5 27.8 41.3 46.9 25.3 25.3 42.1 42.3 33.1 33.6 / / /
GPT-2+BiLSTM-CRF 35.5 36.0 46.2 52.4 28.7 28.8 47.8 47.9 38.6 38.9 / / /
TNT-KID 33.6 33.8 46.0 53.6 31.0 32.0 48.5 48.5 35.9 36.1 / / /
JointKPE 41.7 34.4 35.2 35.0 36.0 29.2 / / / / 36.4 39.1 33.8
Diff-KPE 41.7 34.3 32.3 35.0 35.0 31.4 / / / / 37.8 38.5 32.7
Our approaches - validation set training
DeBERTa 29.2±1.3 29.4±1.3 44.8±0.6 53.9±1.1 26.8±1.1 27.5±1.0 41.9±0.7 41.9±0.7 35.2±0.9 35.4±0.9 27.6±1.5 36.5±0.9 37.2±0.8
DeBERTa MoE 31.1±1.0 31.5±1.1 44.1±0.7 53.7±0.5 26.9±0.3 27.4±0.5 42.5±0.4 42.5±0.4 36.4±2.4 36.8±2.6 27.7±1.0 38.0±0.7 38.9±0.6
DeBERTa RNN 31.2±1.0 31.5±1.0 47.7±0.7 56.6±0.8 29.2±1.5 29.6±1.5 44.1±0.7 44.2±0.7 36.2±1.2 36.6±1.3 26.8±1.1 37.2±0.5 38.1±0.4
DeBERTa MoE RNN 32.7±1.1 33.0±1.1 47.2±1.1 56.5±1.0 29.5±1.1 30.1±1.0 44.9±0.3 45.1±0.3 37.4±0.5 37.8±0.6 27.1±0.9 38.9±0.5 40.1±0.4
Our approaches - full training
DeBERTa 34.7±0.6 35.1±0.6 48.1±0.8 57.4±0.8 31.8±1.9 32.1±1.9 54.5±0.6 54.5±0.6 37.5±1.4 37.7±1.4 33.7±0.9 41.4±1.2 41.7±1.2
DeBERTa MoE 33.5±0.7 33.9±0.6 47.1±0.5 56.3±1.1 30.0±0.8 30.2±0.8 55.0±0.4 55.1±0.4 38.7±0.9 39.1±1.0 34.2±0.4 42.4±0.4 42.8±0.4
DeBERTa RNN 34.4±0.6 34.7±0.6 47.6±0.6 57.5±0.8 32.3±1.5 32.6±1.4 53.5±0.1 53.5±0.1 38.1±0.9 38.2±1.0 31.9±2.3 39.8±1.0 40.2±0.8
DeBERTa MoE RNN 36.0±0.3 36.5±0.3 49.4±1.0 58.8±0.4 34.3±0.9 34.7±0.9 54.8±1.0 55.0±0.9 39.0±1.7 39.3±1.8 33.3±1.5 42.5±0.6 42.9±0.5
Table 2: Empirical evaluation of the keyword extractors. We report the average over five runs (with five different random seeds) and the standard deviation. Bold represents the best score according to a specific criterion on a specific dataset.
5 Results and Analysis
In Table 2, we present the results achieved by our approach and compare them to several baselines (see Section 2.2 for details). TF-IDF (we list results from Martinc et al. (2022)), TextRank (we list results from Martinc et al. (2022)), ChatGPT (we report results for the gpt-3.5-turbo version tested in Luo et al. (2023)) and ChatGLM2-6B (we list results from Luo et al. (2023)) algorithms are unsupervised and do not require any training. For the supervised baselines, CopyRNN, CatSeqD, GPT-2, GPT-2+BiLSTM-CRF, and TNT-KID, we list results reported by Martinc et al. (2022), for Diff-KPE, we list results reported by Luo et al. (2023) and for JointKPE, we list results from Sun et al. (2021). Besides testing our custom sequence labelling head (containing MoE, RNN, or both), we also report results for the unmodified pre-trained DeBERTa model with a standard feed-forward token classification head. For it, we apply the same fine-tuning regime as for our experiments with a custom classification head.
In the validation set training scenario, by adding the MoE on top of the DeBERTa model (see configuration DeBERTa MoE), we improve the performance of the vanilla model (see configuration DeBERTa) with a feed-forward token classification head on four out of six datasets according to all evaluation criteria. The improvement is not observed for the two smallest datasets, Krapivin and Inspec. On the other hand, adding the additional RNN head to the vanilla DeBERTa model improves the performance of the base model on all datasets (see configuration DeBERTa RNN) and according to almost all criteria (the only exception being the F1@1 criterion on the openKP dataset). The results therefore indicate a synergy between the MoE and RNN layers, since the model with both MoE and RNN layers manages to substantially improve the base model and also outperforms the DeBERTa MoE and DeBERTa RNN models on four out of six datasets according to all criteria.
In the full training scenario, the DeBERTa MoE model outperforms the vanilla DeBERTa on three out of six datasets according to all evaluation criteria. Interestingly, the DeBERTa MoE model fails on all computer science datasets, which suggests that expert specialisation might be harder on this domain. On the other hand, the DeBERTa MoE RNN model outperforms the vanilla one on all datasets and according to almost all criteria, the only exception again being the F1@1 criterion on the openKP dataset.
In this scenario we also manage to substantially improve the SOTA on the Inspec, KPTimes, and JPTimes datasets according to both criteria. While we outperform the best-performing model JoinKPE in F1@10 on the KP20k and the Krapivin datasets, our best-performing configuration (DeBERTa MoE RNN) on the other hand achieves uncompetitive performance in terms of F1@5 on these two datasets. On the openKP, our best approach (DeBERTa MoE RNN) fails to beat JointKPE and Diff-KPE approaches according to the F1@1 criterion. The performance comparison is nevertheless favourable according to the other two criteria. Finally, the results also show that the unsupervised models in general achieve much worse performance than the supervised models.
5.1 Examining the Specialisation of Experts
Next, we assess how individual experts correlate with the tokens they operate on to determine their specific types of specialisation. During inference on the test sets, we extracted the expert with the highest weight (i.e., the expert contributing the most out of the top $k$) for each token in the input sequence. We then correlated the selected expert per token with the token’s grammatical and semantic attributes. On the grammatical/lexical level, we consider: whether a token is a specific word, part-of-speech tag (POS), punctuation character or not (Punct), or stopword or not (Stop.). On the semantic level, we examine the predicted labels for each token (Labels), whether a given token is a named entity (bin. NE), and if so, the specific type of named entity (NE). This way, we try to identify different types of specialisation, i.e., whether specialisation is rooted in syntax or semantics.
We annotated the tokens with POS tags and NE labels using the SpaCy library Honnibal et al. (2020). In total, we obtain six different categories of specialisation, which we annotate on the dataset level (i.e., we obtain annotations for the concatenation of documents belonging to a specific dataset). Since we work with categorical values, we measure the correlation between the highest-weighted experts and each category by employing Cramér’s V statistic Cramér (1999) with the bias correction proposed in Bergsma (2013). We consider the two training regimes, validation set and full training, separately, to examine the impact of data size. We present the results in Table 3.
Experts to Inspec Krapivin KP20k KPTimes JPTimes openKP
Validation set training
Words 0.609 0.456 0.592 0.465 0.525 0.475
Punct. 0.701 0.391 0.688 0.076 0.063 0.000
Stop. 0.614 0.403 0.555 0.160 0.096 0.339
POS 0.571 0.376 0.528 0.260 0.323 0.305
Labels 0.257 0.184 0.135 0.093 0.107 0.153
bin. NE 0.085 0.025 0.056 0.271 0.198 0.148
NE 0.070 0.045 0.046 0.246 0.286 0.162
Full training
Words 0.575 0.468 0.459 0.465 0.367 0.521
Punct. 0.128 0.081 0.048 0.082 0.040 0.009
Stop. 0.257 0.105 0.115 0.168 0.092 0.224
POS 0.415 0.265 0.262 0.267 0.165 0.255
Labels 0.228 0.120 0.341 0.184 0.089 0.269
bin. NE 0.039 0.029 0.043 0.238 0.131 0.171
NE 0.048 0.028 0.052 0.237 0.139 0.167
Table 3: Results of the Cramer’s V statistic for correlation between the most weighted expert for a specific token and the specialisation categories examined, across Validation set training and Full training regimes.
In the validation set training scenario, a strong correlation between specific words and specific experts is observed on all datasets. This is not surprising, since MoE works on the token level. On the other hand, the correlation between experts and labels (I, O, and B) tends to be moderate at best (i.e., between 0.1 and 0.3) on most datasets. One can also observe a moderate (and on some datasets strong, i.e., more than 0.5) correlation between specific experts and POS tags, suggesting that specialisation is to a large extent influenced by the syntax of the text.
By looking at the other three types of correlation, namely the correlation between experts and NEs, between experts and stopword tokens, and between experts and punctuation tokens, one can see a very distinct difference between the computer science and other datasets. While there is a strong tendency for experts to specialize for stopword and punctuation tokens on all the computer science datasets (Inspec, Krapivin, and KP20k), this tendency is much weaker on other (news and web page) datasets, where one can observe almost no correlation. On the contrary, here one can observe a specialisation of experts for NEs, which is almost non-existent on the computer science datasets. A correlation is almost the same if we binarize the NE sequence (i.e., to NE/non-NE tokens) or if we look for correlations between experts and 23 labels returned by the SpaCy NE recognition pipeline. The specialisation for NEs can be perhaps explained by the fact that news datasets contain much more NEs than scientific papers and that many of them are in fact labelled as keywords. The same is true for the OpenKP dataset that includes web pages sampled from the index of Bing search engine content-oriented pages, i.e., news pages, multi-media pages from video sites, or indexing pages. The strong correlation with simple syntactic patterns on the computer science datasets can be on the other hand explained by the lack of obvious semantic clues, such as NEs, in these datasets.
When comparing the specialisation in the validation set training scenario and the full training scenario, one can see that correlation to labels, words, NEs, and binary NEs is mostly the same. On the other hand, one can see a substantial decrease in correlation to POS tags, stopwords and punctuation on the computer science datasets, when the models are trained on more data. The assumption is that this is the consequence of expert specialisation becoming more complex when the amount of data allows it.
6 Conclusion
In summary, we investigated the potential of MoE for keyword extraction tasks. We introduced SEKE, a novel architecture that combines MoE and RNN built upon the DeBERTa model for English. This approach achieved superior performance compared to existing SOTA models on standard datasets for keyword extraction. Further analysis confirmed the benefits of MoE layers during fine-tuning, especially with adequate data. The study revealed that adding recurrent layers can help mitigate the data limitations, showing the feasibility of the approach even on small training datasets with less than 2,000 documents. Nevertheless, the specialization analysis shows that the size of the training data still influences the specialisation and that more data allows for more complex specialisation patterns to develop. In the future, we will try to pinpoint the exact dataset size thresholds, down to which the proposed approach is still feasible.
The analysis also revealed that specialisation is domain-specific. Current results indicate that in text with higher complexity (i.e., scientific papers) experts to a larger extent rely on syntactical patterns than in the less complex text, where experts to a larger extent specialize for processing of semantic clues. This hypothesis will be further tested in the future by applying the approach on new domains and languages other than English. This expansion will also further test the generalisability of the proposed approach, reducing the language and domain limitations of the current study.
Another limitations is the usage of an BiLSTM encoder in the custom token classification head, which makes the inference and fine-tuning slightly slower and more computationally demanding than in the vanilla DeBERTa model. While the difference in time was negligible in all our experiments, it might have a visible effect, if the model is applied on a very large dataset. We should also point out the reliance on SpaCY in the specialisation analysis, which is a general domain NLP tool, rarely employed on specialized texts, such as scientific articles. Its usage might have affected the reliability of the analysis’ correlation results by extracting fewer or incorrect NEs.
Finally, we can observe that despite the impressive improvements in the development of foundational LLMs that offer impressive zero-shot capabilities, keyword extraction task can still not be sufficiently solved by employing these models, which lag substantially behind most supervised approaches on most datasets. The biggest advantage of supervised approaches is their ability to adapt to the specifics of the syntax, semantics, content, genre and keyword tagging regime of the specific corpus Meng et al. (2017). Since this cannot be done in a zero-shot setting, in the future we will try to adapt LLMs to the specific corpora and keyword extraction regimes with fine-tuning and in-context learning.
7 Acknowledgements
The authors acknowledge the financial support from the Slovenian Research Agency for research core funding (No. P2-0103) and projects Embeddings-based techniques for Media Monitoring Applications (EMMA, No. L2-50070) and Large Language Models for Digital Humanistics (LLM4DH, No. GC-0002). A Young Researcher Grant PR-12394 supported the work of the last author.
Appendix A Appendix: Examples of keyword identification
This section presents examples of extracted keywords on randomly selected instances from all datasets.
1.) Inspec:
The Bagsik Oscillator without complex numbers. We argue that the analysis of the so-called Bagsik Oscillator, recently published by Piotrowski and Sladkowski (2001), is erroneous due to: (1) the incorrect banking data used and (2) the application of statistical mechanism apparatus to processes that are totally deterministic.
Predicted keywords: incorrect banking data, statistical mechanism apparatus, bagsik oscillator
True keywords: data, statistical mechanism apparatus, complex numbers, bagsik oscillator
2.) Krapivin:
Solving Equations in the Relational Algebra. Enumerating all solutions of a relational algebra equation is a natural and powerful operation which, when added as a query language primitive to the nested relational algebra, yields a query language for nested relational databases, equivalent to the well-known powerset algebra. We study sparse equations, which are equations with at most polynomially many solutions. We look at their complexity and compare their expressive power with that of similar notions in the powerset algebra.
True keywords: relational algebra, nested relation, equation
Predicted keywords: equations, relational algebra, powerset
3.) KP20k:
An algebraic approach to guarantee harmonic balance method using grobner base. Harmonic balance (HB) method is well known principle for analyzing periodic oscillations on nonlinear networks and systems. Because the HB method has a truncation error, approximated solutions have been guaranteed by error bounds. However, its numerical computation is very time-consuming compared with solving the HB equation. This paper proposes an algebraic representation of the error bound using Grobner base. The algebraic representation enables to decrease the computational cost of the error bound considerably. Moreover, using singular points of the algebraic representation, we can obtain accurate break points of the error bound by collisions.
True keywords: algebraic representation, singular point, error bound, grobner base, harmonic balance method
Predicted keywords: grobner base, bound, harmonic balance, harmonic balance method
4.) KPTimes:
Afghan Police Chief Is Killed as He Tries to Turn Tide Against Taliban. A hard-charging Afghan police chief with deep experience in Afghanistan’s long conflict with the Taliban was killed in a blast on Sunday in the country’s eastern Nangarhar Province, which has been under threat from the Taliban and affiliates of the Islamic State. The police chief, Gen. Zarawar Zahid, was visiting an outpost in the Hisarak district when explosives placed near the outpost detonated, according to Attaullah Khogyani, a spokesman for the governor of Nangarhar. One of General Zahid’s bodyguards was wounded, Mr. Khogyani said. More Than 14 Years After U.S. Invasion, the Taliban Control Large Parts of Afghanistan At least one-fifth of the country is controlled or contested by the Taliban. The Taliban claimed responsibility for the killing, according to a statement by the insurgency’s spokesman, Zabiullah Mujahid. The attack came a week after twin bombings outside Afghanistan’s Ministry of Defense killed at least 40 people , including several senior security officials. Nangarhar, which borders Pakistan, has faced mounting security perils over the past couple of years, with new Islamic State affiliates complicating the threat from the Taliban. Zabihullah Zmarai, a member of the provincial council, said the Islamic State posed a danger in five districts, despite repeated operations by the Afghan Army. Out of the 22 districts, only six are secure, he said. The Taliban’s presence across nearly a dozen districts varies, Mr. Zmarai said. But the Hisarak district faced a collapse in recent weeks. That drew the attention of General Zahid, who had gone there to supervise a counterattack. Over the past decade, he rose from a bodyguard to a well-regarded police chief of several volatile provinces. His postings included two stints as the police chief of southeastern Ghazni Province, and one term each in Zabul and Paktika Provinces. General Zahid was seen as a hands-on commander, often arriving at the front lines unannounced. When a major cultural event drew world leaders to the ancient city of Ghazni, the general was photographed riding around the city on the back of a motorcycle to check on security measures. He had been wounded twice and had lost two brothers during the decades of war in Afghanistan. In June, he took part in clashes with Pakistani forces that erupted on the border. In a Facebook video that he posted, he appeared beside two mortars and shouted to his men, "Strike hard enough to blow up Nawaz Sharif’s home," referring to the prime minister of Pakistan. Sediq Sediqqi, the spokesman for Afghanistan’s Ministry of Interior Affairs, called General Zahid "one of the bravest commanders of Afghan police.cHe lost his life on the front line of duty in the fight against terrorism," Mr. Sediqqi said.
True keywords: afghanistan, taliban, zarawar zahid, bombs, nangarhar province,terrorism
Predicted keywords: afghanistan, zarawar zahid, taliban
5.) JPTimes:
Photo report: FOODEX Japan 2013. FoodEx is the largest trade exhibition for food and drinks in Asia, with about 70,000 visitors checking out the products presented by hundreds of participating companies. I was lucky to enter as press; otherwise, visitors must be affiliated with the food industry and pay to enter. The FoodEx menu is global, including everything from cherry beer from Germany and premium Mexican tequila to top-class French and Chinese dumplings. The event was a rare chance to try out both well-known and exotic foods and even see professionals making them. In addition to booths offering traditional Japanese favorites such as udon and maguro sashimi, there were plenty of innovative twists, such as dorayaki , a sweet snack made of two pancakes and a red-bean filling, that came in coffee and tomato flavors. While I was there I was lucky to catch the World Sushi Cup Japan 2013, where top chefs from around the world were competing and presenting a wide range of styles that you would not normally see in Japan, like the flower makizushi above.
True keywords: foodex
Predicted keywords: foodex, japan, food
6.) openKP:
"Quickly create pointandclick games and virtual tours for Windows native PSP iPhone and iPod Touch web apps No programming required very easy to use Free edition contains all the main features Includes free drawing tool and music composer NEW Create Games for iPhone DOWNLOAD FREE GAMES CREATED WITH ADVENTURE MAKER WATCH THE GUIDED TOUR Whats New in version 45 Whats New in version 44
True keywords: virtual tours, adventure maker, no programming required
Predicted keywords: virtual tours